OpenStudy (anonymous):

7r- 7 = 2r + 18 Answer this problem correctly and get a medal!

OpenStudy (anonymous):

r = 5 7r-7=2r+18 -2r -2r ------------- 5r-7=+18 +7 +7 ------------- 5r=25/5 r=5